# Structured Probabilistic Coding

### 21 Dec 2023 v1

@@ Introduction
We conduct experiments on various classification and regression tasks, as shown in Table 1. Concretely, following Barbieri et al. (2020), we experiment on 7 classification tasks about tweet analysis on social media, i.e., EmojiEval (Barbieri et al. 2018), EmotionEval (Mohammad et al. 2018), HatEval (Basile et al. 2019), IronyEval (Hee, Lefever, and Hoste 2018), OffensEval (Zampieri et al. 2019), SentiEval (Rosenthal et al. 2014), and StanceEval (Mohammad et al. 2016). To better evaluate the generalization of the method for cross-domain scenes, we also experiment on 3 emotion-related datasets from different domains, i.e., ISEAR (Scherer and Wallbott 1994), MELD (Poria et al. 2019), and GoEmotions (Demszky et al. 2020). Besides, we experiment on 2 regression benchmarks, i.e., STS-B (Cer et al. 2017) and CLAIRE (Roth, Anthonio, and Sauer 2022). See the appendix for more descriptions of datasets and tasks.
@@
We compare against the 4 universal models (i.e., SVM, FastText, BiLSTM, and GPT-3.5-turbo) and 7 representative deep representation learning technologies (i.e., CE/MSE, CE/MSE+CP, CE+AT, CE+SCL, VIB, MINE-IB, and MEIB). VIB, MINE-IB, and MEIB belong to probabilistic embedding methods, while the others belong to deterministic embedding methods. For these representation learning technologies, we use pre-trained language models, i.e., BERT (Devlin et al. 2019), and RoBERTa (Liu et al. 2019), as the model backbone for fine-tuning on downstream tasks. Concretely, we use bert-base-uncased [^2] https://huggingface.co/ and roberta-base2 to initialize BERT and RoBERTa for fine-tuning on downstream tasks, respectively.
SVM (Cortes and Vapnik 1995) is a machine learning algorithm with a hinge loss that aims to find the best hyperplane to separate data points into different classes. FastText (Joulin et al. 2017) is an efficient text classification method with negative log-likelihood loss based on n-gram features and a hierarchical softmax. BiLSTM is a bidirectional recurrent neural network (Hochreiter and Schmidhuber 1997) that can be used for classification with cross-entropy loss. GPT-3.5-turbo [^3] https://chat.openai.com is an enhanced generative pre-trained transformer model based on text-davinci-003, optimized for chatting. [^4] We present the zero-shot results of the snapshot from June 13th 2023 based on specific inputs, including task descriptions, task instructions, and evaluation texts.
CE/MSE means a fine-tuned baseline with a cross-entropy (CE) loss for classification tasks or a mean squared error (MSE) loss for regression tasks. CE/MSE+CP (Pereyra et al. 2017) is an entropy regularization method that fits a deterministic network by optimizing an objective that combines the CE/MSE loss with a confidence penalty term. CE+AT (Miyato, Dai, and Goodfellow 2017) uses a cross-entropy objective with classical adversarial training. CE+SCL (Gunel et al. 2021) combines CE and supervised contrastive learning (SCL) (Khosla et al. 2020). SCL allows for multiple positives per anchor, thus adapting contrastive learning to the fully supervised setting. VIB (Alemi et al. 2017; Mahabadi, Belinkov, and Henderson 2021) is an efficient variational estimation method based of the information bottleneck (IB) principle (Tishby and Zaslavsky 2015). MINE-IB (Belghazi et al. 2018) is an neural estimation method of the IB principle with a continuous setting. MEIB (An, Jammalamadaka, and Chong 2023) is a variational approach to stochastic embedding in which maximum conditional entropy acts as the bottleneck. MEIB encourages obvious inputs that can be easily classified to take broader embedding areas by assigning larger entropy.
@@
All experiments are conducted on a single NVIDIA Tesla A100 80GB card. The validation sets are used to tune hyperparameters and choose the optimal model. For each method, we run five random seeds and report the average result of the test sets. The network parameters are optimized by using Adamax optimizer (Kingma and Ba 2015) with the learning rate of $5e^{-5}$, the weight decay term of $\{0,0.01,0.001\}$. The dropout rate is set to 0.2. The trade-off parameter $\beta$ is searched from $\{0.001,0.01,0.1,1,10\}$, and the parameter $\gamma$ is searched from $\{0.01,0.05,0.1,0.5,1,5,10\}$. Besides, we conduct experiments using an epoch number of $20$, a total batch size of $128$, and a maximum token length of $128$. The maximum patience for early stopping is set to 5 epochs. More experimental details are listed in the Appendix.
@@
The overall results for 10 classification tasks are summarized in Table 2. Our SPC consistently obtains the best average performance over comparison methods. When using BERT and RoBERTa backbones, SPC can enhance average performance by +3.1% and +4.0% compared to CE for all classification tasks, respectively. The results indicate the good generalization ability of our method to unseen test sets and show the superiority of our method on classification tasks. We notice that SPC achieves big improvements for HatEval and IronyEval, i.e., +14.3% macro-F1 scores and +7.3% F1 scores of ironic, respectively. In HateEval, there is a topic distribution disparity between the validation set and the test set. Additionally, the IronyEval task requires complex semantic understanding, with subtle differences between ironic and non-ironic texts. These results indicate that our SPC has a good generalization capability on the above specific scenarios, i.e., topic shifts and subtle semantic labels.
@@
Table 3 presents the overall results of comparison methods in terms of Spearman and Pearson correlation coefficients for two regression tasks. It shows that SPC obtains better regression results on both datasets. Besides, when using RoBERTa backbone, compared to MSE, SPC achieves +1.5% absolute improvements in terms of the average performance. This demonstrates the superiority and generalization of SPC to unseen test sets on regression tasks.
Methods EmotionEval $\rightarrow$ GoEmotions MELD $\rightarrow$ GoEmotions ISEAR $\rightarrow$ GoEmotions Avg. CE 73.79$\pm$2.57 30.71$\pm$0.54 42.99$\pm$2.10 49.16 CE+AT 72.54$\pm$3.89 32.05$\pm$1.69 44.11$\pm$1.44 49.57 VIB 74.73$\pm$3.52 30.50$\pm$1.05 41.88$\pm$1.65 49.03 MEIB 75.55$\pm$2.05 30.11$\pm$1.33 42.10$\pm$0.61 49.25 SPC 77.47$\pm$2.46 33.95$\pm$1.16 44.36$\pm$1.29 51.93
@@
We conduct ablation studies by removing the structured regularization (w/o Structure) and probabilistic coding (w/o Probabilistic). For classification, Table 4 shows the ablation results on all tasks. When removing the structured regularization, SPC w/o Structure obtains inferior performance in terms of all classification metrics. When further removing probabilistic embedding w/o Probabilistic, the results decline significantly. It reveals the effectiveness of structured regularization and probabilistic coding. Since the label space of regression is a one-dimensional real number, our SPC is degraded to probabilistic coding. The ablation w/o Probabilistic MSE is equivalent to the standard MSE. From Table 3, the average performance declines 1.5% on regression metrics, which confirms the effectiveness of probabilistic coding for regression.
@@
We choose emotion-related benchmarks including EmotionEval, MELD, ISEAR, and GoEmotions, which aim to predict the emotional state but are collected from different domains. To implement OOD scenarios, we train the model on the original training set from a source domain, select the best model based on the validation set of the source domain, and test on the test set of a target domain. To avoid the interference of label mapping bias between different taxonomies, each model is trained on the dataset with coarse-grained taxonomy to predict the label for another dataset with fine-grained taxonomy. Table 5 shows the performance under OOD scenarios. Our SPC obtains the best results on all OOD settings. The fact exhibits SPC’s better generalization capabilities in handling OOD scenarios across different domain shifts. On the one hand, SPC leverages variational inference in the output space, which can better control and utilize randomness and uncertainty of data. On the other hand, SPC introduces the structure information of target task space, making latent space probability distribution better reflect task-related information and generalizing the model to new data.
@@
Methods Noisy EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg. CE 10% 30.66$\pm$0.89 78.15$\pm$0.88 47.06$\pm$5.40 56.90$\pm$4.58 79.46$\pm$0.80 72.36$\pm$0.74 67.39$\pm$1.86 70.40$\pm$0.97 42.01$\pm$1.94 47.85$\pm$1.08 59.22 VIB 10% 30.74$\pm$0.48 77.78$\pm$2.05 47.64$\pm$1.57 58.66$\pm$10.60 79.96$\pm$0.73 72.13$\pm$0.54 67.54$\pm$1.20 70.85$\pm$0.33 38.63$\pm$0.89 47.30$\pm$1.65 59.12 MEIB 10% 31.02$\pm$0.47 78.94$\pm$0.46 49.28$\pm$4.58 57.21$\pm$8.07 80.19$\pm$0.83 72.09$\pm$0.68 68.26$\pm$0.68 70.85$\pm$0.38 38.67$\pm$0.97 46.93$\pm$1.06 59.34 SPC 10% 32.25$\pm$0.69 78.88$\pm$0.47 56.13$\pm$5.36 58.88$\pm$4.94 80.14$\pm$0.28 72.76$\pm$0.06 68.57$\pm$1.01 71.10$\pm$0.62 43.90$\pm$1.13 49.32$\pm$1.22 61.19 CE 20% 31.96$\pm$0.88 77.01$\pm$1.51 49.12$\pm$0.72 60.82$\pm$3.56 79.54$\pm$1.64 72.06$\pm$0.63 68.49$\pm$1.20 70.32$\pm$0.26 40.16$\pm$1.94 47.78$\pm$0.84 59.73 VIB 20% 30.46$\pm$0.59 79.00$\pm$0.49 47.91$\pm$2.20 60.67$\pm$4.82 79.15$\pm$1.22 72.26$\pm$0.29 66.83$\pm$0.52 71.02$\pm$0.25 39.33$\pm$1.47 47.83$\pm$1.38 59.45 MEIB 20% 30.84$\pm$0.75 78.38$\pm$0.88 50.02$\pm$5.18 55.12$\pm$7.07 78.17$\pm$2.55 71.63$\pm$1.11 68.05$\pm$0.81 70.68$\pm$0.38 39.09$\pm$0.87 47.29$\pm$1.22 58.93 SPC 20% 32.51$\pm$0.83 77.97$\pm$1.12 55.41$\pm$6.00 66.40$\pm$4.26 80.33$\pm$0.48 72.50$\pm$0.55 68.89$\pm$1.60 71.10$\pm$0.39 43.96$\pm$0.50 49.04$\pm$0.42 61.81 CE 30% 31.82$\pm$0.75 77.61$\pm$0.90 50.69$\pm$2.80 58.90$\pm$11.45 78.11$\pm$2.07 70.15$\pm$0.5 69.07$\pm$1.07 70.74$\pm$0.56 40.61$\pm$2.06 47.76$\pm$2.29 59.55 VIB 30% 30.85$\pm$0.53 78.23$\pm$0.79 48.22$\pm$1.97 58.81$\pm$8.84 79.38$\pm$0.62 72.15$\pm$0.52 67.59$\pm$0.93 70.27$\pm$0.74 38.71$\pm$1.19 47.16$\pm$1.32 59.14 MEIB 30% 30.74$\pm$0.87 77.99$\pm$0.69 49.98$\pm$4.00 57.57$\pm$5.19 72.53$\pm$5.53 71.83$\pm$0.40 67.88$\pm$0.68 69.86$\pm$1.24 39.39$\pm$1.06 47.43$\pm$1.52 58.52 SPC 30% 32.27$\pm$0.48 78.13$\pm$1.13 56.04$\pm$7.44 59.27$\pm$8.56 80.32$\pm$0.53 72.44$\pm$0.36 69.77$\pm$0.93 70.91$\pm$0.30 43.29$\pm$0.53 49.82$\pm$2.55 61.23
@@
To assess the quality of the representations, we evaluate the clustering performance of output representations obtained by different optimization objectives. Following Hu et al. (2023), we apply silhouette coefficient (SC) and adjusted rand index (ARI) to measure the clustering ability relevant to input data and target labels, respectively. Figure 3 shows SC and ARI of representations learned by various learning objectives. According to the results, SPC achieves higher ARI or SC values compared to other objectives (CE, VIB, and MEIB) across most datasets. It suggests that SPC effectively achieves a balance between data compression and task encoding, thereby promoting the generalization of pre-trained language models for downstream tasks.
@@
According to the nature of embeddings, representation learning approaches can be broadly categorized into deterministic embedding and probabilistic embedding.

### 23 Feb 2024 v4

@@ Introduction
We conduct experiments on various classification and regression tasks. Concretely, following Barbieri et al. (2020), we experiment on 7 classification tasks about tweet analysis on social media, i.e., EmojiEval (Barbieri et al. 2018), EmotionEval (Mohammad et al. 2018), HatEval (Basile et al. 2019), IronyEval (Hee, Lefever, and Hoste 2018), OffensEval (Zampieri et al. 2019), SentiEval (Rosenthal, Farra, and Nakov 2017), and StanceEval (Mohammad et al. 2016). To better evaluate the generalization of the method for cross-domain scenes, we also experiment on 3 emotion-related datasets from different domains, i.e., ISEAR (Scherer and Wallbott 1994), MELD (Poria et al. 2019), and GoEmotions (Demszky et al. 2020). Besides, we experiment on 2 regression benchmarks, i.e., STS-B (Cer et al. 2017) and CLAIRE (Roth, Anthonio, and Sauer 2022). See the appendix for more descriptions of datasets and tasks.
@@
We compare against the 4 universal models (i.e., SVM, FastText, BiLSTM, and GPT-3.5) and 7 representative deep representation learning technologies (i.e., CE/MSE, CE+CP, CE/MSE+AT, CE+SCL, VIB, MINE-IB, and MEIB). VIB, MINE-IB, and MEIB belong to probabilistic embedding methods, while the others typically belong to deterministic embedding methods. For these representation learning technologies, we use pre-trained language models, i.e., BERT (Devlin et al. 2019), and RoBERTa (Liu et al. 2019), as the backbone models for fine-tuning on downstream tasks. Concretely, we use bert-base-uncased [^2] https://huggingface.co/ and roberta-base2 to initialize BERT and RoBERTa for fine-tuning on downstream tasks, respectively.
SVM (Cortes and Vapnik 1995) is a machine learning algorithm with a hinge loss that aims to find the best hyperplane to separate data points into different classes. FastText (Joulin et al. 2017) is an efficient text classification method with negative log-likelihood loss based on n-gram features and a hierarchical softmax. BiLSTM is a bidirectional recurrent neural network (Hochreiter and Schmidhuber 1997) that can be used for classification with cross-entropy loss. GPT-3.5333https://chat.openai.com is an enhanced generative pre-trained transformer model based on text-davinci-003, optimized for chatting. [^4] We present the zero-shot results of the GPT-3.5-turbo snapshot from June 13th 2023 based on specific inputs, including task descriptions, task instructions, and evaluation texts.
CE/MSE means a fine-tuned baseline with a cross-entropy (CE) loss for classification tasks or a mean squared error (MSE) loss for regression tasks. CE+CP (Pereyra et al. 2017) is an entropy regularization method that fits a deterministic network by optimizing an objective that combines the CE loss with a confidence penalty term. CE/MSE+AT (Miyato, Dai, and Goodfellow 2017) uses CE/MSE with classical adversarial training. CE+SCL (Gunel et al. 2021) combines CE and supervised contrastive learning (SCL) (Khosla et al. 2020). SCL allows for multiple positives per anchor, thus adapting contrastive learning to the fully supervised setting. VIB (Alemi et al. 2017; Mahabadi, Belinkov, and Henderson 2021) is an efficient variational estimation method of the information bottleneck (IB) principle (Tishby and Zaslavsky 2015). MINE-IB (Belghazi et al. 2018) is an neural estimation method of the IB principle with a continuous setting. MEIB (An, Jammalamadaka, and Chong 2023) is a variational approach to stochastic embedding in which maximum conditional entropy acts as the bottleneck. MEIB encourages obvious inputs that can be easily classified to take broader embedding areas by assigning larger entropy.
@@
All experiments are conducted on a single NVIDIA Tesla A100 80GB card. The validation sets are used to tune hyperparameters and choose the optimal model. For each method, we run five random seeds and report the average result of the test sets. The network parameters are optimized by using Adamax optimizer (Kingma and Ba 2015) with the learning rate of $5e^{-5}$, the weight decay coefficient of $\{0,0.01,0.001\}$. The dropout rate is set to 0.2. The trade-off parameter $\beta$ is searched from $\{0.001,0.01,0.1,1,10\}$, and the parameter $\gamma$ is searched from $\{0.01,0.05,0.1,0.5,1,5,10\}$. More experimental details are listed in the Appendix.
Methods EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg.
${}^{\dagger}$ 29.30 64.70 36.70 61.70 52.30 62.90 67.30 - - - -
${}^{\dagger}$ 25.80 65.20 50.60 63.10 73.40 62.90 65.40 - - - -
${}^{\dagger}$ 24.70 66.00 52.60 62.80 71.70 58.30 59.40 - - - -
GPT-3.5 6.34$\pm$0.01 73.23$\pm$0.18 48.30$\pm$0.11 66.81$\pm$3.26 63.71$\pm$0.13 40.40$\pm$3.13 39.45$\pm$0.10 67.22$\pm$0.09 41.46$\pm$0.11 25.21$\pm$0.08 47.21
BERT backbone
CE 22.30$\pm$0.60 76.05$\pm$1.41 44.67$\pm$1.78 59.38$\pm$3.01 80.16$\pm$1.26 70.54$\pm$0.44 65.21$\pm$0.71 67.17$\pm$0.78 39.80$\pm$0.84 46.29$\pm$0.79 57.16
CE+CP 21.91$\pm$0.71 76.28$\pm$1.20 45.97$\pm$2.93 64.06$\pm$2.41 78.99$\pm$1.57 70.68$\pm$0.31 65.83$\pm$0.39 67.20$\pm$0.95 39.54$\pm$1.61 46.39$\pm$0.63 57.69
CE+AT 22.93$\pm$0.70 75.08$\pm$1.23 46.30$\pm$3.61 64.23$\pm$2.04 79.68$\pm$1.59 70.55$\pm$0.57 66.46$\pm$1.13 65.70$\pm$0.69 39.84$\pm$0.38 47.37$\pm$0.54 57.81
CE+SCL 21.72$\pm$0.51 75.43$\pm$1.37 45.86$\pm$1.15 65.39$\pm$2.46 80.20$\pm$0.56 70.70$\pm$0.79 65.34$\pm$0.60 67.54$\pm$0.64 40.00$\pm$1.96 46.50$\pm$0.46 57.87
VIB 21.31$\pm$0.62 77.37$\pm$0.71 45.99$\pm$1.93 63.82$\pm$1.00 80.37$\pm$1.11 70.39$\pm$0.31 65.43$\pm$0.60 67.24$\pm$0.57 38.52$\pm$0.51 45.89$\pm$1.10 57.63
MINE-IB 21.29$\pm$0.31 76.60$\pm$0.41 47.64$\pm$2.11 65.86$\pm$2.57 78.67$\pm$2.28 69.85$\pm$0.54 65.35$\pm$0.88 67.62$\pm$0.40 41.23$\pm$0.67 46.87$\pm$0.42 58.10
MEIB 21.87$\pm$0.73 76.70$\pm$0.82 48.27$\pm$1.72 65.87$\pm$2.14 80.49$\pm$0.81 70.55$\pm$0.57 65.59$\pm$1.58 67.44$\pm$0.50 39.30$\pm$0.61 46.26$\pm$0.81 58.23
SPC 24.19$\pm$1.55 77.15$\pm$0.73 57.48$\pm$2.99 65.85$\pm$1.07 80.65$\pm$0.78 70.74$\pm$0.12 67.17$\pm$1.08 68.94$\pm$0.35 42.68$\pm$0.94 47.62$\pm$1.38 60.25
RoBERTa backbone
CE 30.25$\pm$1.32 77.41$\pm$1.33 45.49$\pm$4.70 57.99$\pm$4.96 78.74$\pm$2.20 71.80$\pm$0.93 66.78$\pm$1.34 70.00$\pm$0.45 39.23$\pm$0.41 46.64$\pm$1.15 58.43
CE+CP 31.12$\pm$0.84 77.54$\pm$0.70 48.59$\pm$3.28 58.75$\pm$6.19 79.50$\pm$0.98 72.82$\pm$0.29 66.89$\pm$1.67 70.58$\pm$0.71 40.74$\pm$0.89 47.98$\pm$0.65 59.45
CE+AT 32.00$\pm$0.93 77.30$\pm$1.07 44.71$\pm$4.76 60.17$\pm$3.17 79.81$\pm$1.11 72.51$\pm$0.44 67.81$\pm$0.95 70.97$\pm$0.68 40.10$\pm$0.60 47.89$\pm$1.21 59.33
CE+SCL 31.09$\pm$1.85 76.98$\pm$2.02 49.51$\pm$2.86 60.71$\pm$4.23 80.39$\pm$0.83 73.16$\pm$0.44 66.73$\pm$1.54 70.26$\pm$0.45 40.64$\pm$1.02 47.87$\pm$0.86 59.72
VIB 29.71$\pm$0.79 77.99$\pm$0.86 49.39$\pm$3.08 59.93$\pm$4.57 79.63$\pm$0.66 72.81$\pm$0.39 68.40$\pm$0.52 70.74$\pm$0.44 38.94$\pm$0.55 46.23$\pm$0.18 59.38
MINE-IB 31.70$\pm$0.45 78.79$\pm$0.58 46.39$\pm$2.82 57.39$\pm$8.27 79.76$\pm$0.67 72.85$\pm$0.56 67.27$\pm$1.00 70.15$\pm$0.58 41.80$\pm$2.14 48.88$\pm$1.04 59.50
MEIB 29.94$\pm$1.30 78.73$\pm$0.90 49.34$\pm$2.42 60.54$\pm$2.70 79.68$\pm$0.98 72.78$\pm$0.29 67.89$\pm$1.70 70.86$\pm$0.61 39.00$\pm$0.37 47.18$\pm$1.15 59.59
SPC $*{}^{*}$$\pm$0.48 79.01$\pm$0.61 $*{}^{*}$$\pm$1.32 65.31$\pm$1.91 80.98$\pm$1.36 72.96$\pm$0.22 $*{}^{*}$$\pm$0.63 $*{}^{*}$$\pm$0.59 $*{}^{*}$$\pm$0.29 $*{}^{*}$$\pm$0.60 62.47
Table 1: Classification evaluation (%) on 10 benchmark datasets. ${}^{\dagger}$ means the results are from Barbieri et al. (2020). For other methods, we run five random seeds and report the average result on test sets. BERT and RoBERTa are the backbone models for deep representation learning technologies. Best results for each dataset are highlighted in bold. $*{}^{*}$ represents statistical significance over state-of-the-art scores under the $t$ test ($p<0.05$).
Methods EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg.
SPC 32.54$\pm$0.48 79.01$\pm$0.61 59.80$\pm$1.32 65.31$\pm$1.91 80.98$\pm$1.36 72.96$\pm$0.22 69.02$\pm$0.63 71.01$\pm$0.59 43.99$\pm$0.29 50.04$\pm$0.60 62.47
  - w/o S. 30.98$\pm$0.89 78.60$\pm$0.54 56.64$\pm$8.75 62.12$\pm$6.97 79.16$\pm$1.28 72.23$\pm$0.77 68.90$\pm$0.60 70.79$\pm$0.22 43.75$\pm$0.67 46.61$\pm$1.26 60.98
  - w/o S. & P. 30.25$\pm$1.32 77.41$\pm$1.33 45.49$\pm$4.70 57.99$\pm$4.96 78.74$\pm$2.20 71.80$\pm$0.93 66.78$\pm$1.34 70.00$\pm$0.45 39.23$\pm$0.41 46.64$\pm$1.15 58.43
Table 2: Ablation results (%) on classification tasks. w/o S. indicates removing the structured regularization. w/o P. refers to removing probabilistic coding. We experiment with RoBERTa backbone.
Methods STS-B CLAIRE Avg.
Spearman Pearson Spearman Pearson
MSE 88.33$\pm$0.32 88.80$\pm$0.36 50.37$\pm$5.90 49.10$\pm$5.74 69.15
MSE+AT 88.40$\pm$0.50 89.01$\pm$0.37 53.09$\pm$0.64 51.87$\pm$0.65 70.59
VIB 88.45$\pm$0.50 89.01$\pm$0.40 52.86$\pm$0.88 51.66$\pm$0.78 70.49
MEIB 88.61$\pm$0.14 89.13$\pm$0.17 52.85$\pm$0.72 51.39$\pm$0.81 70.50
SPC 88.71$\pm$0.19 89.31$\pm$0.24 53.11$\pm$0.95 52.21$\pm$0.81 70.84
Table 3: Regression evaluation (%) on 2 benchmark datasets with RoBERTa backbone. For each method, we run five random seeds and report the average result on test sets.
@@
The overall results for 10 classification tasks are summarized in Table 1. Our SPC consistently obtains the best average performance over comparison methods. When using BERT and RoBERTa backbones, SPC can enhance average performance by +3.1% and +4.0% compared to CE for all classification tasks, respectively. The results indicate the good generalization ability of our method to unseen test sets and show the superiority on classification tasks. We notice that SPC achieves big improvements for HatEval and IronyEval, i.e., +14.3% macro-F1 scores and +7.3% F1 scores of the ironic class, respectively. In HateEval, there is a topic distribution disparity between the validation set and the test set. Additionally, the IronyEval task requires complex semantic understanding, with subtle differences between ironic and non-ironic texts. These results indicate that our SPC has a good generalization capability on the above specific scenarios, i.e., topic shifts and subtle semantic labels.
@@
Table 3 presents the overall results of comparison methods in terms of Spearman and Pearson correlation coefficients for two regression tasks. SPC obtains better regression results on both datasets. Besides, when using RoBERTa backbone, compared to MSE, SPC achieves +1.7% absolute improvements in terms of the average performance. This demonstrates the superiority and generalization of SPC to unseen test sets on regression tasks.
@@
We conduct ablation studies by removing the structured regularization (w/o S.) and probabilistic coding (w/o P.). For classification, Table 2 shows the ablation results on all tasks. When removing the structured regularization, the ablated model obtains inferior performance in terms of all classification metrics. When further removing probabilistic coding, the results decline significantly. It reveals the effectiveness of structured regularization and probabilistic coding. For regression, since the label space is a one-dimensional real-valued score, our SPC is degraded to probabilistic coding. The ablation removing probabilistic coding is equivalent to MSE. From Table 3, the average performance declines 1.7% on regression metrics, which confirms the effectiveness of probabilistic coding for regression.
@@
Methods EmotionEval $\rightarrow$ GoEmotions ISEAR $\rightarrow$ GoEmotions MELD $\rightarrow$ GoEmotions Avg.
CE 73.79$\pm$2.57 42.99$\pm$2.10 30.71$\pm$0.54 49.16
CE+AT 72.54$\pm$3.89 44.11$\pm$1.44 32.05$\pm$1.69 49.57
VIB 74.73$\pm$3.52 41.88$\pm$1.65 30.50$\pm$1.05 49.03
MEIB 75.55$\pm$2.05 42.10$\pm$0.61 30.11$\pm$1.33 49.25
SPC 77.47$\pm$2.46 44.36$\pm$1.29 33.95$\pm$1.16 51.93
@@
We choose emotion-related benchmarks including EmotionEval, ISEAR, MELD, and GoEmotions, which aim to predict the emotional state but are collected from different domains. To implement OOD scenarios, we train the model on the original training set from a source domain, select the best model based on the validation set of the source domain, and test on the test set of a target domain. To avoid the interference of label mapping bias between different taxonomies, each model is trained on the dataset with coarse-grained taxonomy to predict the label for another dataset with fine-grained taxonomy. Table 4 shows the performance under OOD scenarios. Our SPC obtains the best results on all OOD settings. The fact exhibits SPC’s better generalization capabilities in handling OOD scenarios across different domain shifts. It indicates that SPC can better control and utilize randomness and uncertainty of data under the probabilistic coding framework. Moreover, the structure regularization from the target task space makes the learned probabilistic representations better reflect task information and generalize the model to new data.
@@
Methods Noisy EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg.
CE 10% 30.66$\pm$0.89 78.15$\pm$0.88 47.06$\pm$5.40 56.90$\pm$4.58 79.46$\pm$0.80 72.36$\pm$0.74 67.39$\pm$1.86 70.40$\pm$0.97 42.01$\pm$1.94 47.85$\pm$1.08 59.22
VIB 10% 30.74$\pm$0.48 77.78$\pm$2.05 47.64$\pm$1.57 58.66$\pm$10.60 79.96$\pm$0.73 72.13$\pm$0.54 67.54$\pm$1.20 70.85$\pm$0.33 38.63$\pm$0.89 47.30$\pm$1.65 59.12
MEIB 10% 31.02$\pm$0.47 78.94$\pm$0.46 49.28$\pm$4.58 57.21$\pm$8.07 80.19$\pm$0.83 72.09$\pm$0.68 68.26$\pm$0.68 70.85$\pm$0.38 38.67$\pm$0.97 46.93$\pm$1.06 59.34
SPC 10% 32.25$\pm$0.69 78.88$\pm$0.47 56.13$\pm$5.36 58.88$\pm$4.94 80.14$\pm$0.28 72.76$\pm$0.06 68.57$\pm$1.01 71.10$\pm$0.62 43.90$\pm$1.13 49.32$\pm$1.22 61.19
CE 20% 31.96$\pm$0.88 77.01$\pm$1.51 49.12$\pm$0.72 60.82$\pm$3.56 79.54$\pm$1.64 72.06$\pm$0.63 68.49$\pm$1.20 70.32$\pm$0.26 40.16$\pm$1.94 47.78$\pm$0.84 59.73
VIB 20% 30.46$\pm$0.59 79.00$\pm$0.49 47.91$\pm$2.20 60.67$\pm$4.82 79.15$\pm$1.22 72.26$\pm$0.29 66.83$\pm$0.52 71.02$\pm$0.25 39.33$\pm$1.47 47.83$\pm$1.38 59.45
MEIB 20% 30.84$\pm$0.75 78.38$\pm$0.88 50.02$\pm$5.18 55.12$\pm$7.07 78.17$\pm$2.55 71.63$\pm$1.11 68.05$\pm$0.81 70.68$\pm$0.38 39.09$\pm$0.87 47.29$\pm$1.22 58.93
SPC 20% 32.51$\pm$0.83 77.97$\pm$1.12 55.41$\pm$6.00 66.40$\pm$4.26 80.33$\pm$0.48 72.50$\pm$0.55 68.89$\pm$1.60 71.10$\pm$0.39 43.96$\pm$0.50 49.04$\pm$0.42 61.81
CE 30% 31.82$\pm$0.75 77.61$\pm$0.90 50.69$\pm$2.80 58.90$\pm$11.45 78.11$\pm$2.07 70.15$\pm$0.5 69.07$\pm$1.07 70.74$\pm$0.56 40.61$\pm$2.06 47.76$\pm$2.29 59.55
VIB 30% 30.85$\pm$0.53 78.23$\pm$0.79 48.22$\pm$1.97 58.81$\pm$8.84 79.38$\pm$0.62 72.15$\pm$0.52 67.59$\pm$0.93 70.27$\pm$0.74 38.71$\pm$1.19 47.16$\pm$1.32 59.14
MEIB 30% 30.74$\pm$0.87 77.99$\pm$0.69 49.98$\pm$4.00 57.57$\pm$5.19 72.53$\pm$5.53 71.83$\pm$0.40 67.88$\pm$0.68 69.86$\pm$1.24 39.39$\pm$1.06 47.43$\pm$1.52 58.52
SPC 30% 32.27$\pm$0.48 78.13$\pm$1.13 56.04$\pm$7.44 59.27$\pm$8.56 80.32$\pm$0.53 72.44$\pm$0.36 69.77$\pm$0.93 70.91$\pm$0.30 43.29$\pm$0.53 49.82$\pm$2.55 61.23
@@
To assess the quality of the representations, we evaluate the clustering performance of output representations obtained by different optimization objectives. Following Hu et al. (2023), we apply silhouette coefficient (SC) and adjusted rand index (ARI) to measure the clustering ability relevant to input data and target labels, respectively. Figure 3 shows SC and ARI of representations learned by various learning objectives. According to the results, SPC achieves higher ARI or SC values compared to other objectives (CE, VIB, and MEIB) across most datasets. It suggests that SPC effectively achieves a balance between data encoding and task prediction, thereby promoting the generalization of pre-trained language models for downstream tasks.
@@
Representation learning provides low-dimensional representations of the data, which can effectively capture the data features and serve for various tasks. According to the nature of embeddings, it can be broadly categorized into deterministic embedding and probabilistic embedding.